### v2

# CURE4Rec: A Benchmark for Recommendation Unlearning with Deeper Influence

Abstract
With increasing privacy concerns in artificial intelligence, regulations have mandated the right to be forgotten, granting individuals the right to withdraw their data from models. Machine unlearning has emerged as a potential solution to enable selective forgetting in models, particularly in recommender systems where historical data contains sensitive user information. Despite recent advances in recommendation unlearning, evaluating unlearning methods comprehensively remains challenging due to the absence of a unified evaluation framework and overlooked aspects of deeper influence, e.g., fairness. To address these gaps, we propose CURE4Rec, the first comprehensive benchmark for recommendation unlearning evaluation. CURE4Rec covers four aspects, i.e., unlearning Completeness, recommendation Utility, unleaRning efficiency, and recommendation fairnEss, under three data selection strategies, i.e., core data, edge data, and random data. Specifically, we consider the deeper influence of unlearning on recommendation fairness and robustness towards data with varying impact levels. We construct multiple datasets with CURE4Rec evaluation and conduct extensive experiments on existing recommendation unlearning methods. Our code is released at https://github.com/xiye7lai/CURE4Rec.
1 Introduction
Over the past few years, growing concerns over information abundance and data leakage have intensified the focus on privacy preservation within artificial intelligence. Regulations such as the General Data Protection Regulation (GDPR) (Union, 2018), the California Consumer Privacy Act (Pardau, 2018)and the Delete Act (Information, 2023) grant individuals the right to be forgotten, requiring the deletion of personal data used in information systems. Nowadays, the ubiquitous application of machine learning models in information systems poses potential risks for memorizing training data (Fredrikson et al., 2015). Consequently, the aforementioned regulations also require forgetting the associated data memory within the trained models, giving rise to the concept of machine unlearning. Recently, machine unlearning has gained increasing popularity in computer vision (Bourtoule et al., 2021; Gupta et al., 2021), natural language processing (Chen & Yang, 2023; Eldan & Russinovich, 2023), and recommender systems (Chen et al., 2022; Li et al., 2023a, b). As recommender systems typically rely on historical interaction data to extract user preferences, the recommendation model inherently contains sensitive user information. Therefore, there is a crucial need for unlearning to preserve privacy. The task of machine unlearning in recommender systems is termed as recommendation unlearning.
While machine unlearning has demonstrated significant potential in preserving user privacy, conducting a comprehensive evaluation of unlearning methods continues to pose difficulties. Various unlearning methods employ distinct evaluation metrics, yet a universally applicable evaluation framework remains absent. Specifically, existing evaluation methods predominantly focus on the unlearning completeness, unlearning efficiency, and its impact on model utility, overlooking the deeper influence of model properties.
In this paper, we identify two overlooked aspects of deeper influence. Firstly, fairness is a crucial consideration for recommendations (Wang et al., 2023), but is often neglected in unlearning evaluations. Ensuring fair recommendation outcomes can avoid user discrimination and enrich the recommendation platform’s understanding of user preferences. Existing studies demonstrate that unlearning can affect the fairness of models (Oesterling et al., 2024). Secondly, existing evaluation methods neglect the influence of various unlearning sets, randomly selecting data for unlearning. Distinct unlearning sets, however, can result in significantly different impacts on model performance (Fan et al., 2024). Performing comprehensive evaluations on different unlearning data contributes to understanding the robustness of unlearning methods.
Figure 1: An illustration of CURE4Rec, a comprehensive benchmark tailored for evaluating recommendation unlearning methods. CURE4Rec evaluates unlearning methods using data with varying levels of unlearning impact on four aspects, i.e., unlearning completeness, recommendation utility, unlearning efficiency, and recommendation fairness.
To address these issues, we introduce CURE4Rec, a comprehensive benchmark specifically designed to evaluate recommendation unlearning methods. As shown in Figure 1, CURE4Rec’s evaluation encompasses four aspects, i.e., unlearning completeness, recommendation utility, unlearning efficiency, and recommendation fairness. Additionally, each aspect is investigated with three data selection strategies, i.e., core data, edge data, and random data. This triadic breakdown tests to reflect the robustness of recommendation unlearning methods towards different unlearning sets. The main contributions of this work are summarized as follows:
  • We introduce CURE4Rec, a comprehensive benchmark tailored for evaluating recommendation unlearning methods. CURE4Rec enables evaluation across multiple aspects, including unlearning completeness, recommendation utility, unlearning efficiency, and recommendation fairness.
  • To the best of our knowledge, we are the first to investigate the impact of unlearning on recommendation fairness, introducing fairness evaluation to comprehensively grasp its impact and proposing additional requirements to consider for further research.
  • We further examine the impact of different unlearning sets. Based on the level of collaboration, we select core data, edge data, and random data to construct unlearning sets respectively, aiming to thoroughly explore the impact towards unlearning completeness, recommendation utility, unlearning efficiency, and recommendation fairness.
  • We offer multiple datasets tailored for evaluation using our CURE4Rec. Furthermore, we conduct extensive experiments across existing recommendation unlearning methods and report their performance (please refer to Figure 2 for an overview of our results).
2 Related Work
2.1 Machine Unlearning
Machine unlearning aims to eliminate the memory of specific data, serving purposes such as privacy protection (Liu et al., 2022) and erasing data biases (Sattigeri et al., 2022; Chen et al., 2024b). According to the level of unlearning completeness, existing machine unlearning methods can be categorized into two approaches, i.e., exact unlearning and approximate unlearning.
Exact Unlearning (EU) aims to completely eliminate the influence of target data on the model. The most straightforward method of exact unlearning is retraining the model from scratch on the updated dataset (removing the target data), but this method incurs a significant computational time cost. To mitigate this cost, existing EU methods revamp the training process via ensemble learning, which limits the retraining cost to sub-datasets or sub-models (Bourtoule et al., 2021; Yan et al., 2022).
Approximate Unlearning (AU) achieves unlearning through direct parameter manipulation, avoiding the significant time cost of retraining. Most AU methods utilize gradients or influence function to estimate the influence of target data and subsequently remove it from models (Sekhari et al., 2021; Wu et al., 2022; Mehta et al., 2022). Alternatively, other methods directly prune or dampen model parameters to achieve unlearning (Wang et al., 2022; Foster et al., 2024).
2.2 Recommendation Unlearning
Recommendation unlearning aims to eliminate the influence of target data within the recommender system (Li et al., 2024b). A naive approach to achieve recommendation unlearning is through the direct application of the classic unlearning method, i.e., SISA (Bourtoule et al., 2021). Due to the collaborative characteristics of recommendation data, tailored methods have been proposed to improve SISA for recommendation unlearning, e.g., RecEraser (Chen et al., 2022) and UltraRE (Li et al., 2023a). In addition to EU methods mentioned above, AU method also enters the scene, utilizing refined influence functions to enable recommendation unlearning (Li et al., 2023b; Zhang et al., 2023). Note that this paper focuses on investigating the model-agnostic approaches. Other approaches focus on specific models, e.g., sequential recommendation (Ye & Lu, 2023), session-based recommendation (Xin et al., 2024), and large language model-based recommendation (Wang et al., 2024; Hu et al., 2024). Note that this paper focuses solely on recommendation unlearning of training data. An alternative line of research, known as attribute unlearning, explores the unlearning of latent user attributes in recommender systems. (Ganhör et al., 2022; Li et al., 2023c; Chen et al., 2024a).
2.3 Machine Unlearning Benchmarks
Emerging research has pioneered early investigation into unlearning benchmarks, focusing on image classification (Choi & Na, 2023), large language models (Maini et al., 2024; Li et al., 2024a; Jin et al., 2024), and diffusion models (Zhang et al., 2024). By proposing new datasets or modifying existing ones, these investigations design depth evaluation metrics within their corresponding domains. However, these benchmarks leave unexplored deeper influence of unlearning on model properties, i.e., fairness and robustness. This exploration is crucial for recommender systems, as alternations in the performance of recommendation models immediately affect recommendation lists, eventually influencing use experience. To the best of our knowledge, we are the first to introduce a recommendation unlearning benchmark, and comprehensively explore the deeper influence of unlearning on recommendation fairness and robustness.
3 CURE4Rec
In this section, we first recall the process of recommendation unlearning, outlining the necessary inputs for evaluations. Then, we introduce evaluation aspects of our proposed CURE4Rec, detailing specific metrics for each aspect. Finally, we present the strategy for unlearning set selection.
3.1 Recommendation Unlearning
The entire process of recommendation unlearning consists of four stages: I) completing learning process to generate the original model; II) determining the unlearning set, i.e., the unlearning target, which is a subset of training data; III) conducting unlearning process based on the original model to produce the unlearned model; and IV) evaluating the unlearned model. To ensure reliable evaluation, we evaluate unlearning methods using identical training and testing data, employing the same learning process to generate the same original model. This ensures that all unlearning methods start from the same baseline in stage I. To investigate unlearning robustness, we select three types of unlearning sets in stage II (Section 3.3). In stage IV, CURE4Rec’s evaluation includes the four aspects (Section 3.2).
In the context of recommendation, unlearning targets may vary among users, items, and user-item interactions. Commonly, recommendation unlearning scenarios focus on user-wise unlearning (Li et al., 2023a). Thus, our benchmark primarily investigates the user-wise unlearning scenarios.
3.2 Evaluation Aspects
Unlearning Completeness.
Unlearning completeness stands as the primary goal and fundamental requirement of recommendation unlearning. Exact unlearning methods inherently guarantee unlearning completeness by retraining, which is the only authorized approach (Thudi et al., 2022). On the other hand, approximate unlearning methods, lacking the ability to achieve authorized unlearning, often require the demonstration of unlearning completeness through theoretical proofs or empirical studies. Therefore, following the completeness evaluation of approximate unlearning in previous studies (Graves et al., 2021; Ma et al., 2022; Li et al., 2023b; Kurmanji et al., 2024), we evaluate unlearning completeness of recommendation unlearning based on the attacking performance of Membership Inference Oracle (MIO).
MIO follows the standard membership inference procedure to evaluate unlearning completeness in image classification task (Graves et al., 2021; Ma et al., 2022). In the context of recommendation, we concatenate user embeddings with the average item embeddings of their respective interacted items as the data features, and the probability of being in the training set as the data label. Please refer to Section 4.5 for more training details. To evaluate unlearning completeness, we query MIO with the unlearned data points. Ideally, MIO outputs 1 (indicating presence in the training set) for the original model and 0 (indicating absence from the training set) for the unlearned model. Since exact unlearning methods guarantee complete unlearning, we only evaluate the completeness of approximate unlearning methods.
Recommendation Utility.
Recommendation unlearning aims to erase the memory of target data within recommender systems without causing harm to the knowledge acquired from the remaining data. Thus, preserving the recommendation utility of the remaining data is another important goal of unlearning. To investigate the impact of unlearning on model utility, we employ two widely used metrics, i.e., Normalized Discounted Cumulative Gain (NDCG) and Hit Ratio (HR), to evaluate the recommendation performance of the unlearned model on the testing set. For both metrics, we truncate the ranked list to 20 items.
Unlearning Efficiency.
Retraining from scratch represents the gold standard in unlearning, but its practical implementation carries a prohibitive computational overhead. Recommender systems encompass hundreds of thousands of users, generating a large amount of unlearning requests. Therefore, improving unlearning efficiency is a crucial goal of recommendation unlearning. We measure unlearning efficiency by the total runtime of the entire unlearning process, i.e., stage III. Note that we enable parallel training for exact unlearning.
Recommendation Fairness.
Previous research has demonstrated that unlearning affects deeper model properties such as fairness (Oesterling et al., 2024). Mitigating the negative impact of unlearning is also an important requirement of unlearning. In this paper, we evaluate the performance fairness of recommendation unlearning from the following two perspectives: i) the fairness between active and inactive groups (A-IGF), and ii) the fairness among different shards (shardGF), as exact unlearning methods divide the datasets into multiple shards.
For A-IGF, we follow the representative user-oriented group fairness research in recommendation (Li et al., 2021). Based on the number of interactions, we classify the top 5% of users as the active group and the remaining 95% users as the inactive group. Active and inactive users are selected outside the unlearning set, because we aim to investigate the impact on the remainder users. Then we compute the difference of the average recommendation utility, i.e., NDCG@20, between active and inactive groups to represent A-IGF. For shardGF, we report the variance of recommendation utility among all shards to compare the shard-level fairness (Rastegarpanah et al., 2019). Note that we do not compute shardGF for approximate unlearning, because these methods do not involve sharding.
3.3 Unlearning Set Selection
Existing evaluation methods typically select data randomly for the unlearning set. However, previous studies have shown that i) poisoned data can be constructed to make it hard to unlearn (Marchant et al., 2022), and ii) different data points have varying difficulty of unlearning (Fan et al., 2024). Motivated by these findings, in this paper, we explore the impact of using varying unlearning sets, which can also reflect the robustness of unlearning.
To significantly demonstrate this impact, we adopt a model-agnostic selection strategy to create three types of unlearning sets: core data (which impacts many other data points), edge data (with minimal impact on others), and random data. Specifically, we regard the user-item interactions as a non-weighted bipartite graph, where users and items are represented as nodes, and an edge connects them if there is an interaction. Existing research suggests that a node’s importance correlates strongly with its centrality in a graph (Haveliwala, 2002; Li et al., 2012; Park et al., 2019). In the context of recommendation, centrality is associated with collaborations, manifested as neighbors in a graph. Thus, we define the importance of a node $x$ as follows:
$I(x)=c(x)\cdot\frac{\sum_{y\in N(x)}c(y)}{\left|N(x)\right|},$ (1)
where $c(x)$ denotes the centrality of node $x$, and $N(x)$ denotes the number of neighbors of node $x$. Due to the collaborative characteristic of recommendation data, we use the degree of node, i.e., the number of first-order neighbors, to compute centrality. Finally, we rank all nodes based on $I(x)$ to select the core data and edge data.
4 Experimental Setup
4.1 Datasets
We conduct experiments on three real-world datasets widely used in recommendation. ML-100K [^1] https://grouplens.org/datasets/movielens/: The MovieLens dataset is one of the most extensively utilized datasets in recommender system research. MovieLens 100k contains 100 thousand individual ratings. ML-1M: MovieLens 1M contains 1 million ratings. ADM [^2] http://jmcauley.ucsd.edu/data/amazon/: The Amazon dataset comprises multiple subsets categorized according to different types of Amazon products. One of these subsets, known as the Amazon Digital Music (ADM) dataset, includes ratings of digital music. Following the widely used pre-processing procedure (He et al., 2017; Wang et al., 2019; He et al., 2020), we convert ratings into implicit feedback. The statistical details of these datasets are summarized in Table 1. To avoid extreme sparsity, we filter out the users and items that have less than 5 interactions. For each dataset, we randomly select 80% ratings as the training set, 10% ratings as the validation set, and the remaining as the test set. The unlearning ratio, i.e., the percentage of unlearning set within the training set, is initially set as 5%. We also explore this ratio within a range of (5%, 10%, 15%, 20%) in Appendix A.4.
Table 1: Summary of datasets.
Dataset User # Item # Interactions # Sparsity
ML-100K 943 1,349 99,287 92.195%
ML-1M 6,040 3,416 999,611 95.155%
ADM 478,235 266,414 836,006 99.999%
4.2 Recommendation Models
Aligning with existing studies on recommendation unlearning (Chen et al., 2022; Li et al., 2023b, a), we use three representative recommendation models based on collaborative filtering for evaluation:
  • WMF: Weighted Matrix Factorization(WMF) (Chen et al., 2020) is a non-sampling recommendation model that treats all missing interactions as negative interactions and assigns them with uniform weights.
  • BPR: Bayesian Personalized Ranking (Rendle et al., 2012) is a widely used recommendation model that uses a Bayesian personalized ranking objective function to optimize matrix factorization.
  • LightGCN: LightGCN (He et al., 2020) is the state-of-the-art collaborative filtering model, which improves recommendation performance by simplifying graph neural networks.
4.3 Unlearning Methods
We consider the following recommendation unlearning methods, including both EU and AU approaches (note that we set the number of shards to 10 for EU and explore other values in Section 5.5):
  • Retrain: Retraining from scratch is the goal standard unlearning method.
  • SISA: SISA (Bourtoule et al., 2021) stands as the classic algorithm for machine unlearning, adaptable to various scenarios, including recommender systems.
  • RecEraser: RecEraser (Chen et al., 2022) is specifically designed for recommendation unlearning, which modifies SISA to boost performance in recommendation tasks.
  • UltraRE: UltraRE (Li et al., 2023a) enhances RecEraser for recommendation tasks by modifying two key stages, i.e., division and aggregation.
  • SCIF: SCIF (Li et al., 2023b) is the first approximate unlearning method in recommendation systems, employing influence functions tailored for recommendation tasks.
4.4 Parameters Settings
In the training phase of original models, we randomly sample 4 negative items for each observed interaction following (He et al., 2017). In the case of model-specific hyper-parameters, we tune them in the ranges suggested by their original papers. In detail, the batch size is set to 512, the learning rate is set to 0.01, the embedding size is set to 32. The maximum number of epochs is set to 500. The early stopping strategy is adopted in our experiments, which terminates the training when NDCG@20 on the validation set does not increase for 5 successive epochs.
4.5 MIO Training Details
Following (Li et al., 2023b), we adopt an ideal concept, i.e., Membership Inference Oracle (MIO), to evaluate unlearning completeness. Specifically, We implement an approximated MIO via a basic three-layer (64, 16, 4) neural network with ReLu and Softmax as activation functions for hidden layers and the output layer respectively. We train the MIO via stochastic gradient descent with 100 epochs and a learning rate of 0.001. The MIO outputs the probability of the queried data point being in the training set. To evaluate the unlearning completeness, we query MIO with the unlearned data points. Ideally, MIO outputs 1 (being in the training set) for the original model while outputs 0 (not being in the training set) for the unlearned model.
4.6 Hardware Information
We run all experiments on the same Ubuntu 20.04 LTS System server with 48-core CPU, 256GB RAM and NVIDIA GeForce RTX 3090 GPU.
Figure 2: A visualized evaluation overview of recommendation unlearning methods in four aspects ($\uparrow$), where the result is the normalized average outcome obtained across all models and datasets, using random data as the unlearning set. The recommendation fairness is measured by A-IGF (fairness between active and inactive users). The higher values represent better performance.
5 Results and Discussion
In this section, we report and analyze the results regarding four evaluation aspects under three selections of unlearning sets. We present a visualized overview of compared recommendation unlearning methods in Figure 2. We observe that apart from unlearning completeness, the AU method (SCIF) demonstrates a significant advantage over EU methods (SISA, RecEraser, and UltraRE), particularly in terms of unlearning efficiency and recommendation fairness. However, it is essential to highlight that unlearning completeness is the primary goal of unlearning. EU methods inherently achieve the highest level of completeness, whereas SCIF can only achieve weak unlearning.
Table 2: Results in terms of unlearning completeness (MIO accuracy - approaching 0.5), recommendation utility (NDCG and HR $\uparrow$), and recommendation fairness (A-IGF - approaching Retrain) for the approximate recommendation unlearning method, where Learn denotes the results before unlearning. Core, random, and edge respectively refer to the selection of the unlearning sets as core data, random data, and edge data.
ML-100K ML-1M ADM
NDCG@20 HR@20 MIO A-IGF NDCG@20 HR@20 MIO A-IGF NDCG@20 HR@20 MIO A-IGF
Learn 0.3215 0.3415 0.722 -0.0450 0.2144 0.2112 0.741 -0.042 0.0277 0.0578 0.756 0.0167
Retrain Core 0.3187 0.3295 0.540 -0.0184 0.2196 0.2174 0.544 -0.0188 0.0221 0.0446 0.555 0.0053
Random 0.2872 0.3353 0.538 -0.0403 0.2124 0.2108 0.547 -0.0507 0.0252 0.0519 0.556 0.0141
Edge 0.3091 0.3140 0.536 -0.0430 0.2148 0.2051 0.546 -0.0518 0.0272 0.0554 0.556 0.0164
SCIF Core 0.2483 0.2382 0.561 -0.0322 0.1865 0.1629 0.569 -0.0213 0.0194 0.0398 0.571 0.0094
Random 0.2699 0.2617 0.563 -0.0268 0.1922 0.1785 0.571 -0.0311 0.0227 0.0461 0.575 0.0106
Edge 0.2894 0.3012 0.601 -0.0375 0.2031 0.1811 0.623 -0.0191 0.0245 0.0502 0.579 0.0103
5.1 Unlearning Completeness
To evaluate the completeness of AU methods, we report the accuracy of MIO in Table 2, where the recommendation model is WMF. Due to the space limit, we report the results of other models in Appendix A.2. Compared the result of SCIF with the performance before unlearning and Retrain after unlearning, we observe that i) both SCIF and Retrain significantly decrease the MIO accuracy, indicating their effectiveness in unlearning; ii) although not significant, there is still a marginal gap between SCIF and Retrain (ground truth), i.e., 4.1% higher accuracy than Retrain on average; and iii) SCIF particularly performance worse on edge data compared to other data types. This discrepancy may be attributed to imprecise influence estimation for this specific data category.
Table 3: Results in terms of recommendation utility for exact recommendation unlearning methods.
ML-100K Retrain SISA RecEraser UltraRE
Core Random Edge Core Random Edge Core Random Edge Core Random Edge
WMF NDCG@20 0.3187 0.2872 0.3091 0.2096 0.2092 0.2041 0.2285 0.2208 0.2109 0.2303 0.2354 0.2149
HR@20 0.3295 0.3353 0.3140 0.2094 0.2049 0.1892 0.2218 0.2142 0.1979 0.2267 0.2282 0.2027
BPR NDCG@20 0.3111 0.3003 0.3043 0.2244 0.2324 0.2298 0.2614 0.2615 0.2694 0.2708 0.2764 0.2743
HR@20 0.3151 0.3028 0.2987 0.2203 0.2259 0.2179 0.2724 0.2658 0.2620 0.2851 0.2813 0.2695
LightGCN NDCG@20 0.3175 0.3121 0.3101 0.1802 0.1932 0.1964 0.2856 0.2905 0.2886 0.2952 0.3069 0.3063
HR@20 0.3250 0.3253 0.3244 0.1724 0.1907 0.1911 0.3053 0.3099 0.3121 0.3123 0.3201 0.3185
ML-1M Retrain SISA RecEraser UltraRE
Core Random Edge Core Random Edge Core Random Edge Core Random Edge
WMF NDCG@20 0.2196 0.2124 0.2148 0.1780 0.1639 0.1714 0.1894 0.1796 0.1838 0.1926 0.1891 0.1970
HR@20 0.2174 0.2108 0.2051 0.1612 0.1485 0.1493 0.1731 0.1592 0.1596 0.1747 0.1680 0.1717
BPR NDCG@20 0.2462 0.2319 0.2336 0.1545 0.1530 0.1628 0.1826 0.1660 0.1860 0.1828 0.1856 0.1913
HR@20 0.2279 0.2162 0.2118 0.1353 0.1329 0.1367 0.1627 0.1450 0.1624 0.1652 0.1632 0.1651
LightGCN NDCG@20 0.2177 0.2108 0.2147 0.1504 0.1533 0.1642 0.1864 0.1863 0.1814 0.1969 0.1867 0.1806
HR@20 0.2138 0.2045 0.2186 0.1365 0.1323 0.1581 0.1825 0.1804 0.1818 0.1907 0.1855 0.1798
ADM Retrain SISA RecEraser UltraRE
Core Random Edge Core Random Edge Core Random Edge Core Random Edge
WMF NDCG@20 0.3691 0.3566 0.3556 0.2720 0.2589 0.2515 0.3373 0.3256 0.3185 0.3420 0.3334 0.3347
HR@20 0.4071 0.3822 0.3848 0.2617 0.2492 0.2471 0.3527 0.3467 0.3203 0.3689 0.3595 0.3501
BPR NDCG@20 0.3566 0.3453 0.3499 0.2806 0.2708 0.2757 0.3286 0.3295 0.3212 0.3325 0.3301 0.3314
HR@20 0.3821 0.3628 0.3718 0.2745 0.2638 0.2611 0.3486 0.3406 0.3483 0.3541 0.3569 0.3608
LightGCN NDCG@20 0.0105 0.0106 0.0096 0.0075 0.0054 0.0048 0.0084 0.0085 0.0079 0.0097 0.0088 0.0086
HR@20 0.0221 0.0234 0.0208 0.0157 0.0112 0.0103 0.0171 0.0176 0.0154 0.0191 0.0185 0.0183
5.2 Recommendation Utility
We report the results in terms of recommendation utility for AU and EU in Tables 2 and 3, respectively. In general, the AU method (SCIF) outperforms the EU methods (SISA, RecEraser, and UltraRE). Employing the same unlearning set, RecEraser and UltraRE consistently outperform SISA across all datasets and models, with UltraRE generally surpassing RecEraser, aligning with previous research (Li et al., 2023a).
For all EU methods, the recommendation utility of unlearning core users is generally higher than that of unlearning random-select or edge users. This is likely due to the removal of data from more interactive users, which typically contains a large amount of ratings. This enables the model to learn more effectively from the smaller amount of remaining training data. Compared with these EU methods, SCIF exhibits the highest recommendation utility, closely resembling that of Retrain. However, SCIF suffers the most substantial performance decline when unlearning core users. This can be attributed to the increased number of interactions involved in calculating the influence function, leading to inaccurate influence estimation that negatively impacts the model utility.
Table 4: Results in terms of unlearning efficiency (running time in seconds $\downarrow$).
Time (s) ML-100K ML-1M ADM
WMF BPR LightGCN WMF BPR LightGCN WMF BPR LightGCN
Retrain Core 4296 5238 4734 7748 9113 8645 3682 6998 5225
Random 4526 5494 5044 8693 9461 10324 3972 7127 5354
Edge 4687 5527 5274 8006 9748 10497 4127 7351 6359
SISA Core 402 488 437 1160 1160 1523 669 1750 1009
Random 467 586 528 1256 1265 1605 717 1842 1246
Edge 442 504 515 1280 1292 1659 751 1902 1077
RecEraser Core 463 582 561 1533 1568 1846 865 1892 1106
Random 476 693 656 1654 1660 1952 912 1945 1490
Edge 489 659 617 1736 1819 1964 965 2032 1190
UltraRE Core 457 591 559 1507 1493 1667 819 1810 1057
Random 482 618 645 1595 1550 1834 901 1862 1283
Edge 466 518 666 1781 1791 1955 923 1904 1368
SCIF Core 289 336 316 784 784 1034 453 1186 682
Random 325 403 368 862 860 1083 497 1242 841
Edge 316 358 359 887 877 1126 520 1282 733
(a) ML-100K
(b) ML-1M
(c) ADM
Figure 3: Results in terms of recommendation fairness for exact recommendation unlearning methods on WMF, where A-IGF (approaching Retrain) and shardGF ($\downarrow$) evaluate the fairness of group-level and shard-level, respectively.
5.3 Unlearning Efficiency
We report the unlearning times in Table 4. In general, SCIF is more efficient than EU methods. Among the EU methods, SISA saves more time compared to RecEraser and UltraRE, because it does not have the complex division and aggregation stage specific to the recommendation scenarios. Due to its design, UltraRE is slightly more efficient than RecEraser. Additionally, EU methods take less time to unlearn core users since they have a larger amount of interaction data. This reduces the amount of data left for retraining. On the contrary, SCIF requires more computations for influence estimation on core users, resulting in higher time costs compared to unlearning random or edge users.
(a) Utility
(b) Efficiency
(c) Group-level Fairness
(d) Shard-level Fairness
Figure 4: Effect of shard number in terms of multiple aspects, i.e., recommendation utility ($\uparrow$), unlearning efficiency ($\downarrow$), group-level fairness (approaching Retrain), and shard-level fairness ($\downarrow$).
5.4 Recommendation Fairness
We also report the recommendation fairness of AU and EU methods in Table 2 and Figure 3, respectively. For the group-level fairness (A-IGF), compared to the AU method (SCIF), EU methods notably worsen unfairness, tending to favor active users. This is primarily attributed to the division stage of EU methods, with this effect becoming more pronounced on larger datasets, i.e., ML-1M and ADM. Moreover, RecEraser and UltraRE, which group active users together instead of randomly, as done by SISA, exacerbate unfairness even further. For the shard-level fairness (shardGF), although to a lesser extent compared to group-level fairness, RecEraser and UltraRE also exacerbate unfairness.
5.5 Effects of Shard Number
We report the effect of shard number in terms of multiple aspects in Figure 4, using WMF on ML-100K. Firstly, as the number of shards increases, the unlearning efficiency improves, but the recommendation utility deteriorates, as confirmed by several previous studies (Chen et al., 2022; Li et al., 2023a). Secondly, the increased shard number further groups the active users into smaller shards, exacerbating the group-level fairness. At the same time, it reduces the discrepancy among all shards, diminishing the shard-level fairness.
6 Conclusion
In this paper, we present a comprehensive benchmark, CURE4Rec, for recommendation unlearning methods, aiming to analyze and inspire further exploration into the deeper influence of recommendation unlearning. Specifically, CURE4Rec covers four evaluation aspects, i.e., unlearning completeness, recommendation utility, unlearning efficiency, and recommendation fairness. Additionally, we investigate unlearning robustness across three unlearning sets, i.e., core data, edge data, and random data. Through extensive experiments, we compare the performance of various recommendation unlearning methods using our proposed benchmark. Our experiments reveal that i) the division-aggregation design of the EU approach has dual implications. On one hand, it inherently achieves unlearning completeness. On the other hand, it compromises other evaluation aspects. and ii) The AU approach, which directly manipulates model parameters, outperforms the EU approach in all aspects except completeness, with less negative influence on model properties, e.g., fairness.
Limitation and Boarder Impact.
This paper proposes a benchmark for recommendation unlearning, comprising four evaluation aspects. This design also offers insights for other unlearning scenarios. Simultaneously, there is considerable room for improvement in the specific evaluation metrics within each aspect. Additionally, the AU approach appears to outperform the EU approach in all aspects except completeness. The trade-off between completeness and other aspects is an intriguing direction that is not discussed in this paper.
Acknowledgments and Disclosure of Funding
This work was supported by the Fundamental Research Funds for the Central Universities 226-2024-00241. We thank all the anonymous reviewers for helpful feedback on early versions of this work.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes] See Section 6.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] See Section 6.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] See Abstract.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See Section 4.4.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [No]
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Section 4.6.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] See Section 4.1.
    2. (b)
      Did you mention the license of the assets? [N/A]
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [No]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [N/A]
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
Appendix A More Results
A.1 Performance Overview
We report a visualized overview of compared recommendation unlearning methods on each dataset in Figure 5. The results are generally consistent with Figure 2.
(a) ML-100K
(b) ML-1M
(c) ADM
Figure 5: A visualized evaluation overview of recommendation unlearning methods in four aspects ($\uparrow$), where the result is the normalized average outcome obtained across all models, using random data as the unlearning set. The recommendation fairness is measured by A-IGF (fairness between active and inactive users).
A.2 Unlearning Completeness
We report the accuracy of MIO in Table 5, where the recommendation model is BPR. We omit the results for LightGCN as we encountered difficulties accurately computing the influence function of SCIF on ML-1M and ADM based on current hardware.
Table 5: Results in terms of unlearning completeness (MIO accuracy - approaching 0.5), recommendation utility (NDCG and HR $\uparrow$), and recommendation fairness (A-IGF - approaching Retrain) for the approximate recommendation unlearning method, where Learn denotes the results before unlearning. Core, random, and edge respectively refer to the selection of the unlearning sets as core data, random data, and edge data.
ML-100K ML-1M ADM
NDCG@20 HR@20 MIO A-IGF NDCG@20 HR@20 MIO A-IGF NDCG@20 HR@20 MIO A-IGF
Learn 0.3195 0.3030 0.724 -0.0246 0.2517 0.2306 0.744 -0.0651 0.0251 0.0510 0.759 0.0194
Retrain Core 0.3111 0.3151 0.536 -0.0217 0.2462 0.2279 0.549 -0.0374 0.0246 0.0504 0.558 0.0066
Random 0.3003 0.3028 0.535 -0.0153 0.2319 0.2162 0.550 -0.0605 0.0203 0.0421 0.561 0.0187
Edge 0.3043 0.2987 0.537 -0.0175 0.2336 0.2118 0.552 -0.0633 0.0203 0.0439 0.555 0.0191
SCIF Core 0.2392 0.2182 0.565 -0.0116 0.1898 0.1636 0.572 -0.0284 0.0171 0.0336 0.573 0.0096
Random 0.2768 0.2824 0.566 -0.0144 0.2159 0.1886 0.576 -0.0372 0.0189 0.0357 0.573 0.0110
Edge 0.2871 0.2905 0.612 -0.0167 0.2231 0.1942 0.635 -0.0481 0.0200 0.0417 0.588 0.0132
A.3 Recommendation Fairness
We report the recommendation fairness of exact unlearning methods on each dataset using BPR and LightGCN recommendation models in Figures 6 and 7, respectively.
We also report the grouping results of active and inactive users after applying three exact unlearning methods, i.e., SISA, RecEraser, UltraRE, on different datasets in Tables 6, 7, and 8. On the one hand, SISA randomly distributes both types of users evenly across groups. On the other hand, RecEraser and UltraRE tend to cluster active users into the same groups, which results in certain groups containing numerous active users while others have almost none. This clustering result explains why RecEraser and UltraRE tend to favor active users, as the concentration of active users in certain groups significantly increases their proportion compared to random distribution, leading to more effective learning but also more severe unfairness.
(a) ML-100K
(b) ML-1M
(c) ADM
Figure 6: Results in terms of recommendation fairness for exact recommendation unlearning methods on BPR, where A-IGF (approaching Retrain) and shardGF ($\downarrow$) evaluate the fairness of group-level and shard-level, respectively.
(a) ML-100K
(b) ML-1M
(c) ADM
Figure 7: Results in terms of recommendation fairness for exact recommendation unlearning methods on LightGCN, where A-IGF (approaching Retrain) and shardGF ($\downarrow$) evaluate the fairness of group-level and shard-level, respectively.
A.4 Unlearning Ratio
We report the effect of unlearning data ratio in terms of multiple aspects in Figure 8, using WMF on ML-100K. We observe consistent results with previous studies (Bourtoule et al., 2021; Chen et al., 2022; Li et al., 2023a). In general, as the ratio of unlearning data increases, the recommendation utility gradually decreases, along with a reduction in the unlearning time. Additionally, a larger unlearning ratio tends to lead to greater fairness.
Table 6: Results of user distribution (active vs. inactive) in each shard on dataset ML-100K. The unlearning data ratio is set to 5%.
ML-100K Group 1 Group 2 Group 3 Group 4 Group 5
Active Inactive Active Inactive Active Inactive Active Inactive Active Inactive
SISA Core 4 86 4 86 3 87 7 83 8 82
Random 4 86 8 82 2 88 3 87 6 84
Edge 3 87 6 84 3 87 6 84 6 84
RecEraser Core 0 90 10 80 0 90 0 90 0 86
Random 0 90 1 89 1 89 0 90 0 90
Edge 0 90 6 84 0 90 0 86 1 89
UltraRE Core 0 89 11 78 0 90 6 83 0 90
Random 0 90 3 86 3 87 1 89 0 89
Edge 0 89 15 75 1 88 2 87 7 83
ML-100K Group 6 Group 7 Group 8 Group 9 Group 10
Active Inactive Active Inactive Active Inactive Active Inactive Active Inactive
SISA Core 1 89 3 86 3 86 6 83 5 84
Random 5 85 4 85 1 88 5 84 6 83
Edge 8 82 2 87 1 88 7 82 2 87
RecEraser Core 0 90 0 90 28 62 0 90 6 84
Random 6 84 9 81 0 90 27 63 0 86
Edge 0 90 9 81 0 90 27 63 1 89
UltraRE Core 7 83 10 80 1 89 0 89 9 81
Random 3 87 1 88 0 89 18 72 15 75
Edge 0 90 1 89 0 90 7 82 11 79
Table 7: Results of user distribution (active vs. inactive) in each shard on dataset ML-1M. The unlearning data ratio is set to 5%.
ML-1M Group 1 Group 2 Group 3 Group 4 Group 5
Active Inactive Active Inactive Active Inactive Active Inactive Active Inactive
SISA Core 28 546 31 543 26 548 30 544 32 542
Random 25 549 34 540 30 544 23 551 35 539
Edge 36 538 20 554 24 550 32 542 27 547
RecEraser Core 44 530 52 522 0 572 20 554 5 569
Random 2 570 79 495 44 530 40 534 5 569
Edge 10 564 41 533 74 500 31 543 0 574
UltraRE Core 0 573 5 569 11 563 12 562 33 541
Random 44 530 6 567 7 567 5 569 13 561
Edge 8 566 9 564 11 563 4 569 23 550
ML-1M group6 Group 7 Group 8 Group 9 Group 10
Active Inactive Active Inactive Active Inactive Active Inactive Active Inactive
SISA Core 23 551 25 549 33 541 28 545 30 543
Random 28 546 38 536 22 552 30 543 21 552
Edge 27 547 39 535 33 541 26 547 22 551
RecEraser Core 64 510 1 573 38 536 35 539 27 547
Random 32 542 1 573 61 513 2 572 20 554
Edge 24 550 2 572 91 483 4 568 9 565
UltraRE Core 44 530 27 547 18 556 32 541 104 470
Random 3 571 0 574 14 559 7 567 187 387
Edge 0 575 49 525 26 548 155 419 1 573
Table 8: Results of user distribution (active vs inactive) in each shard on dataset ADM. The unlearning data ratio is set to 5%.
ADM Group 1 Group 2 Group 3 Group 4 Group 5
Active Inactive Active Inactive Active Inactive Active Inactive Active Inactive
SISA Core 96 2078 113 2061 126 2048 117 2057 106 2068
Random 108 2066 112 2062 95 2079 110 2064 84 2090
Edge 112 2062 105 2069 100 2074 120 2054 106 2068
RecEraser Core 429 1745 0 2174 8 2166 0 2174 0 2169
Random 0 2169 0 2174 453 1721 159 2015 84 2090
Edge 149 2025 84 2090 379 1795 7 2167 0 2169
UltraRE Core 91 2083 160 2013 65 2108 65 2109 88 2086
Random 41 2132 80 2094 361 1813 81 2093 56 2117
Edge 82 2092 11 2162 201 1972 53 2120 330 1844
ADM Group 6 Group 7 Group 8 Group 9 Group 10
Active Inactive Active Inactive Active Inactive Active Inactive Active Inactive
SISA Core 98 2075 112 2061 99 2074 115 2058 104 2069
Random 119 2054 114 2059 135 2038 116 2057 93 2080
Edge 109 2064 106 2067 111 2062 102 2071 115 2058
RecEraser Core 1 2173 97 2077 388 1786 121 2053 42 2132
Random 9 2165 0 2174 0 2174 4 2170 377 1797
Edge 456 1718 0 2174 0 2174 10 2164 1 2173
UltraRE Core 65 2109 173 2001 147 2026 123 2051 109 2063
Random 200 1973 50 2124 137 2036 48 2125 32 2142
Edge 82 2091 58 2116 121 2052 55 2119 93 2081
(a) Utility
(b) Efficiency
(c) Group-level Fairness
(d) Shard-level Fairness
Figure 8: Effect of unlearning ratio in terms of multiple aspects, i.e., recommendation utility ($\uparrow$), unlearning efficiency ($\downarrow$), group-level fairness (approaching Retrain), and shard-level fairness ($\downarrow$).